# Simulating Optical Coherent Nonlinear Response for High-Intensity Excitation

### v1

@@ Appendix A Optimized phase-cycling scheme
    We list signals originating due to aliasing of the required phase-matched signal, as defined by Eq. (6). Specifically, we find all combinations of $\alpha_{j}+IP_{j}$, with integer $I\in[0,N)$. We have chosen the range of $I$ to include all relevant aliased signals. $I=0$ corresponds the phase-matching condition of the required signal.
@@
Appendix B Peak-amplitude calculation
We present the procedure used to calculate the FWM amplitude as function of the intensity of excitation pulses in Sec. III . One can use the peak of the 2D spectrum if the signal-to-noise ratio (SNR) and linewidth are independent of the intensity of the excitation pulses. Both of these conditions are not satisfied for our measurements. The SNR ratio improves and excitation-induced dephasing (EID) Li et al. (2006) leads to increase in the linewidth at higher intensities. Thus, we need to estimate the amplitude while accounting for the above behaviour.
@@ Appendix B Peak-amplitude calculation
EID results in broadening of the exciton peak as the excitation intensity is increased. Consequently, the value of the peak amplitude will result in a lower estimate of the FWM amplitude. We take a cross-diagonal slice across the peak of the noise-corrected spectrum, as shown in Fig. 11. All the values in this slice are summed to obtain a value proportional to total FWM signal intensity; integration along the frequency axis accounts for the effect of increase in the peak linewidth.
@@
The integration of values outside the peak is nearly zero due to noise subtractions, which allows us to correctly estimate the signal strength. A square root of the integrated intensity gives us the value of the FWM signal ampltitude.

### v2

@@ Appendix B Optimized phase-cycling scheme
    We list signals originating due to aliasing of the required phase-matched signal, as defined by Eq. (6). Specifically, we find all combinations of $\alpha_{j}+IP_{j}$, with integer $I\in[0,N)$. We have chosen the range of $I$ to include all relevant aliased signals. $I=0$ corresponds to the phase-matching condition of the required signal.
@@
Appendix C Peak-amplitude calculation
We present the procedure used to calculate the FWM amplitude as function of the intensity of excitation pulses in Sec. III . One can use the peak of the 2D spectrum if the signal-to-noise ratio (SNR) and linewidth are independent of the intensity of the excitation pulses. Both of these conditions are not satisfied for our measurements. The SNR improves and excitation-induced dephasing (EID) Li et al. (2006) leads to increase in the linewidth at higher intensities. Thus, we need to estimate the amplitude while accounting for the above behaviour.
@@ Appendix C Peak-amplitude calculation
EID results in broadening of the exciton peak as the excitation intensity is increased. Consequently, the value of the peak amplitude will result in a lower estimate of the FWM amplitude. We take a cross-diagonal slice across the peak of the noise-corrected spectrum, as shown in Fig. 11. All the values in this slice are summed to obtain a value proportional to total FWM signal intensity; integration along the frequency axis accounts for the effect of increase in the peak linewidth. The integration of values outside the peak is nearly zero due to noise subtractions, which allows us to correctly estimate the signal strength. A square root of the integrated intensity gives us the value of the FWM signal ampltitude.
@@
Appendix D OBEs for diamond system
We present the OBEs for an exciton-biexciton diamond-type system, depicted in Fig. 5. To ensure the cross-linear excitation scheme, we use a sequence of four pulses: the $1^{st}$ and $4^{th}$ are horizontally polarized, and the $2^{nd}$ and $3^{rd}$ are vertically polarized. The total electric field is given as:
$E_{total}(t,r)=E^{H}_{1}+E^{V}_{2}+E^{V}_{3}+E^{H}_{4}$ (15)
where the electric fields $E^{j}_{i}(t,r)$ are defined similarly to those in Eq. 1, with the polarization (H or V) specified by the superscript $j$.
The OBEs for the density matrix elements of the diamond-type system are:
${\dot{\rho}}_{HH}$ $=-\Gamma_{1}\rho_{HH}+\Gamma_{3}\rho_{BB}+\frac{i\mu_{GH}E_{H}}{\hbar}\Big{(}\rho_{GH}-\rho_{HG}\Big{)}$
$\quad+\frac{i\mu_{HB}E_{H}}{\hbar}\Big{(}\rho_{BH}-\rho_{HB}\Big{)},$ (16a)
${\dot{\rho}}_{VV}$ $=-\Gamma_{2}\rho_{VV}+\Gamma_{3}\rho_{BB}+\frac{i\mu_{GV}E_{V}}{\hbar}\Big{(}\rho_{GV}-\rho_{VG}\Big{)}$
$\quad+\frac{i\mu_{VB}E_{V}}{\hbar}\Big{(}\rho_{BV}-\rho_{VB}\Big{)},$ (16b)
${\dot{\rho}}_{BB}$ $=-2\Gamma_{3}\rho_{BB}+\frac{i\mu_{HB}E_{H}}{\hbar}\Big{(}\rho_{HB}-\rho_{BH}\Big{)}$
$\quad+\frac{i\mu_{VB}E_{V}}{\hbar}\Big{(}\rho_{VB}-\rho_{BV}\Big{)},$ (16c)
${\dot{\rho}}_{GH}$ $=(i\omega_{HG}-\gamma_{GH})\rho_{GH}+\frac{i\mu_{GH}E_{H}}{\hbar}\big{(}\rho_{HH}-\rho_{GG}\big{)}$
$\quad+\frac{i}{\hbar}\Big{(}\mu_{GV}E_{V}\rho_{VH}-\mu_{HB}E_{H}\rho_{GB}\Big{)},$ (16d)
${\dot{\rho}}_{GV}$ $=(i\omega_{VG}-\gamma_{GV})\rho_{GV}+\frac{i\mu_{GV}E_{V}}{\hbar}\big{(}\rho_{VV}-\rho_{GG}\big{)}$
$\quad+\frac{i}{\hbar}\Big{(}\mu_{GH}E_{H}\rho_{HV}-\mu_{BV}E_{V}\rho_{GB}\Big{)},$ (16e)
${\dot{\rho}}_{HB}$ $=(i\omega_{BH}-\gamma_{HB})\rho_{HB}+\frac{i\mu_{HB}E_{H}}{\hbar}\big{(}\rho_{BB}-\rho_{HH}\big{)}$
$\quad+\frac{i}{\hbar}\Big{(}\mu_{GH}E_{H}\rho_{GB}-\mu_{VB}E_{V}\rho_{HV}\Big{)},$ (16f)
${\dot{\rho}}_{VB}$ $=(i\omega_{BV}-\gamma_{VB})\rho_{VB}+\frac{i\mu_{VB}E_{V}}{\hbar}\big{(}\rho_{BB}-\rho_{VV}\big{)}$
$\quad+\frac{iE_{H}}{\hbar}\Big{(}\mu_{VG}\rho_{GB}-\mu_{HB}\rho_{VH}\big{)},$ (16g)
${\dot{\rho}}_{HV}$ $=-\gamma_{HV}\rho_{HV}+\frac{i}{\hbar}\Big{(}\mu_{HG}E_{H}\rho_{GV}-\mu_{VB}E_{V}\rho_{HB}\Big{)}$
$\quad+\frac{i}{\hbar}\Big{(}\mu_{HB}E_{H}\rho_{BV}-\mu_{VG}E_{V}\rho_{HG}\Big{)},$ (16h)
${\dot{\rho}}_{GB}$ $=(i\omega_{BG}-\gamma_{GB})\rho_{GB}+\frac{i}{\hbar}E_{H}\Big{(}\mu_{GH}\rho_{HB}-\mu_{HB}\rho_{GH}\Big{)}$
$\quad+\frac{i}{\hbar}E_{V}\Big{(}\mu_{GV}\rho_{VB}-\mu_{VB}\rho_{GV}\Big{)},$ (16i)
where
  • $\rho_{ii}$ : Population in state |$\ket{i}$
  • $\rho_{ij}$ : Coherence between states |$\ket{i}$ and |$\ket{j}$
  • $\mu_{ij}$ : Transition dipole moment for transition between states |$\ket{i}$ and |$\ket{j}$
  • $\Gamma_{i}$ : Population decay rate for state |$\ket{i}$
  • $\gamma_{ij}$ : Coherence decay rate between states |$\ket{i}$ and |$\ket{j}$
  • $\omega_{ij}=\omega_{i}-\omega_{j}$
with, $i\neq j=$ $G$, $H$, $V$, $B$, and $\omega_{GH}=\omega_{GV}$. Similar to Appendix A, we solve these OBEs after applying the RWA.